# Searching for strong lensing by late-type galaxies in UNIONS

### v1

@@
11institutetext: Institute of Physics, Laboratory of Astrophysics, Ecole Polytechnique Fdrale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland
11email: javier.acevedobarroso@epfl.ch
22institutetext: SCITAS, Ecole Polytechnique Fédérale de Lausanne (EPFL), 1015 Lausanne, Switzerland 33institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí i Franquès 1, 08028 Barcelona, Spain 44institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig de Lluís Companys 23, 08010 Barcelona, Spain 55institutetext: Laboratoire dAstrophysique de Marseille, UMR7326, Aix-Marseille Universit, CNRS, CNES, 13013 Marseille, France 66institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS, and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 77institutetext: Oskar Klein Centre, Department of Physics, Stockholm University, SE-106 91 Stockholm, Sweden 88institutetext: University of Applied Sciences and Arts of Northwestern Switzerland, School of Engineering, 5210 Windisch, Switzerland 99institutetext: Institute of Cosmology and Gravitation, University of Portsmouth, Burnaby Rd, Portsmouth PO1 3FX, UK 1010institutetext: Department of Physics and Astronomy, University of British Columbia, Vancouver, BC V6T 1Z1, Canada $1$institutetext: Canadian Astronomy Data Centre, Herzberg Astronomy and Astrophysics, National Research Council, 5071 West Saanich Rd Victoria BC V9E 2E7 1212institutetext: LIRA, Observatoire de Paris, Universit PSL, CNRS, Place Jules Janssen 92195, Meudon, France 1313institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, ON, N2L 3G1, Canada 1414institutetext: Waterloo Centre for Astrophysics, Waterloo, ON, N2L 3G1, Canada 1515institutetext: Perimeter Institute for Theoretical Physics, 31 Caroline St. N., Waterloo, ON, N2L 2Y5, Canada 1616institutetext: Institute for Astronomy, University of Hawaii, 2680 Woodlawn Drive, Honolulu, HI 96822, USA
@@
[…]
(Received September 15, 1996; accepted March 16, 1997)
Recent wide-field galaxy surveys have led to an explosion in numbers of galaxy-scale strong gravitational lens candidates. However, the vast majority feature massive luminous red galaxies as the main deflectors, with late-type galaxies being vastly under-represented. This work presents a dedicated search for lensing by edge-on late-type galaxies in the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). The search covers $3600\,\deg^{2}$ of $r$-band observations taken from the Canada-France-Hawaii Telescope. We consider all sources with magnitudes in the range $20.517<r<$ 20.5, without any colour preselection, yielding a parent sample of seven million sources. We characterise our parent sample via the visual inspection of 120 000120000120\,000120 000 sources selected at random. From it, we estimate, with a 68% confidence interval, that 1 in every 30 0003000030\,00030 000 sources is an edge-on lens candidate, with at least eight high-quality candidates in the parent sample. This corresponds to 1 candidate per 17 0001700017\,00017 000 edge-on late-type galaxies. Our search relies on a convolutional neural network (CNN) to select a reduced sample of candidates, followed by a visual inspection to curate the final sample. The CNN is trained from scratch using simulated $r$-band observations of edge-on lenses, and real observations of non-lenses. We find 61 good edge-on lens candidates using the CNN. Moreover, combining the CNN candidates with those found serendipitously, and those identified while characterising the parent sample, we discovered 4 grade A, 20 grade B, and 58 grade C edge-on lens candidates; effectively doubling the known sample of these systems. We also discovered 16 grade A, 16 grade B, and 18 grade C lens candidates of other types. Finally, based on the characterisation of the parent sample, we estimate that our search found around 60% of the bright grade A and B edge-on lens candidates within the parent sample.
@@ 1 Introduction
Strong gravitational lensing has become an essential tool for probing the extragalactic Universe. Its applications range from the detection and study of high-redshift sources (e.g., Atek et al. 2023; Bradley et al. 2023), including the mapping of the mass distributions responsible for the lensing effect (e.g., Shajib et al. 2019; Etherington et al. 2022; Mahler et al. 2023), to the constraining of cosmological parameters, most notably $H_{0}$ when the lensed source exhibits time variability (e.g., Wong et al. 2020; Birrer et al. 2020; Kelly et al. 2023; Chen et al. 2024). Additionally, the analysis of flux ratios between multiply-lensed images, alongside small-scale distortions in the lensed arcs, allows for the detection of mass clumps in the line of sight (e.g., Vegetti et al. 2010; Nightingale et al. 2023; Nierenberg et al. 2024; Gilman et al. 2024; Nightingale et al. 2024). Furthermore, combining lens modelling with a second probe of mass can break degeneracies in our mass models that limit the individual analyses, such as the baryonic and dark matter degeneracy (e.g., Auger et al. 2010; Schuldt et al. 2019; Sheu et al. 2024), the bulge-disc degeneracy for late-type galaxies (e.g., Dutton et al. 2011; Suyu et al. 2012), or the mass-sheet degeneracy, which affects lens models in general (e.g., Shajib et al. 2023; Khadka et al. 2024). In particular, galaxy-scale lensing, when combined with deep spectroscopic observations or multi-band photometry, can probe the stellar initial mass function (IMF) of the lensing galaxies, see Ferreras et al. (2010); Czoske et al. (2012); Dutton & Treu (2014); Brewer et al. (2014); Sonnenfeld et al. (2019).
@@
However, these studies are limited by their sample size, and their results can change with improved samples and statistics. This is especially the case for late-type galaxies, for which the sample of confirmed lenses is in the few tens. Whether the IMF is the same for late-type galaxies as it is for early-type, in what way the IMF might differ in the bulge from the disc for a given galaxy, and the universality of the IMF for late-type galaxies, are all open questions that can be tackled with large samples of late-type lens galaxies. In this context, the Sloan WFC Edge-on Late-type Lens Survey (SWELLS; WFC wide field camera) launched a campaign in the early 2010 decade to find and study lensing by edge-on late-type galaxies (edge-on lenses), while trying to answer the aforementioned questions. This series of six papers includes: the curation and characterisation of a spectroscopically selected sample of edge-on lenses (Treu et al. 2011); the combination of lens modelling with spatially-resolved kinematics to break the disc-halo degeneracy and probe the IMF for a single galaxy (Dutton et al. 2011; Barnabè et al. 2012); the systematic comparison between the total mass inside the Einstein radius and the stellar mass estimated from stellar population synthesis (SPS) models to study the IMF for medium-sized samples ( $\approx\,20$) of edge-on lenses (Brewer et al. 2012, 2014); and the study of the IMF normalisation of the bulge and the disc independently for a sample of five edge-on lenses using lens modelling, kinematics, and SPS (Dutton et al. 2013). To date, the SWELLS papers represent the largest study of edge-on lenses, but it is still limited to a couple tens of systems.
@@
Moreover, galaxy-scale strong lensing events are very rare. The likelihood of observing these events depends on the mass distribution of the deflectors and the redshift distributions of foreground lenses and background sources. As a result, only about one source in thousands exhibits lensed images (e.g., Oguri & Marshall 2010; Collett 2015; Ferrami & Wyithe 2024; Acevedo Barroso et al. 2024), and those are not easy to see in ground-based imaging, both because of the small image separation on the plane of the sky and because of the low luminosity contrast between the lens light and the background source. This is true when the lens galaxy is a massive luminous red galaxy (LRG), but the situation is even worse for the less massive late-type galaxies. In fact, only about a hundred lens candidates have a late-type deflector, out of approximately twenty thousand lens candidates known by the community (see Vernardos et al., in prep.). More so, even though the number of lens candidates is growing orders of magnitude faster since the adoption of machine learning in lens finding, the number of late-type candidates has been growing at a much more modest rate, hinting at the necessity of lens searches that target them directly.
@@
Recent massive optical lens searches are targeting wide-sky surveys to maximise the volume probed, while also using neural networks to select samples of potential lens candidates. These surveys include the Dark Energy Survey (DES; Jacobs et al. 2019a, b; Rojas et al. 2022; O’Donnell et al. 2022), the Kilo-Degree Survey (KiDS; Petrillo et al. 2017, 2019; He et al. 2020; Li et al. 2020, 2021), the Subaru Hyper Suprime-Cam (HSC; Cañameras et al. 2021; Shu et al. 2022; Jaelani et al. 2024; Schuldt et al. 2025), the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Survey (Huang et al. 2020, 2021; Stein et al. 2022; Storfer et al. 2024), the Euclid Early Release Observations (Nagam et al. 2025), and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Savary et al. 2022). Most of the analyses rely on convolutional neural networks as the main classifier, but visual transformers are also being explored as an alternative (Grespan et al. 2024; González et al. 2025).
@@
Given the extremely low prevalence of lensing events, it is common to restrict lens searches to the most massive galaxies, often making a colour preselection for LRG. These preselections remove most of the late-type galaxy lenses, excluding them from the lens search even before the CNN inspect them. However, even if there is no colour preselection, there are still not enough confirmed lenses to train neural networks using only real observations. Instead, the CNN are trained on mock observations simulating the lensing effect. These mock lenses are being refined, from the first ones by Petrillo et al. (2017), randomly adding fully simulated lensed sources to LRG, to the recent use of real deep HST observations for the potential sources, and the dynamics of the potential deflectors to select a reasonable lensing mass (Cañameras et al. 2020; Rojas et al. 2022; Savary et al. 2022). The problem with this methodology is that it limits the search to the lens candidates that look the most similar to the mock lenses used to train the networks. And, since these studies mostly use LRG for the deflector galaxies, then late-type lenses are also indirectly excluded, even if they are part of the parent sample inspected by the networks.
@@
In this work, we search for galaxy-scale lensing by late-type edge-on disc galaxies (hereafter ‘edge-on lenses’) in the Canada-France Imaging Survey (CFIS), now a part of the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). We choose this survey because of its excellent image quality, large area coverage, and its complete overlap with the on-going Euclid Wide Survey (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024). This represents the second lens search specifically targeting edge-on lenses in optical data, after Sygnet et al. (2010), who also used data from the Canada-France-Hawaii Telescope (CFHT). However, this is the first time it is being done using machine learning and the methodology popularised in the last decade. Our main objective is to substantially increase the sample of known edge-on lens candidates, while also exploring why they are not being found by the recent massive lens searches. Additionally, we aim to estimate the prevalence and total number of edge-on lenses in the data, and develop a framework to evaluate the performance of our CNN-search without relying on simulated lenses.
This paper is organised as follows. We present the UNIONS-CFIS observations and the selection of the parent sample in Sect. 2. Then, in Sect. 3, we characterise the parent sample by estimating the number of lenses in it. We describe the process to simulate mock observations of edge-on lenses in Sect. 4, and present the training and deployment of the CNN in Sect. 5. We present spectroscopic follow-up of two candidates from the Nordic Optical Telescope (NOT) in Sect. 6. Finally, we discuss our findings and conclude in Tables 2 and 8, respectively.
@@ 2 Data
We use r-band data from CFIS data release 3, conducted using the 3.6-m CFHT on Maunakea, Hawaii. CFIS provides deep r- and u-band imaging to UNIONS, a collaboration of wide-field imaging surveys of the northern sky involving CFIS, as well as members of the Panoramic Survey Telescope And Rapid Response System (Pan-STARRS) team, the Waterloo Hawaii Institute for Astronomy G-band Survey (WHIGS) team, and the Wide Imaging with Subaru Hyper Suprime-Cam of the Euclid Sky (WISHES) team. Deep i-band and moderately deep z-band imaging data are provided by Pan-STARRS, deep g-band data are provided by WHIGS, and deep z-band data are provided by the WISHES team. UNIONS aims to provide ground-based complementary data for the Euclid mission, as well as maximising the science output of these ground-based large-area deep imaging surveys (Ibata et al. 2017; Guinot et al. 2022).
@@
We classify detections with flux radius larger than one standard deviation above the median value as extended sources. We then extract postage stamps of $66\times 6$ 66 pixels ($12.256″12.256″$$\times$$$) for every extended source within the range $20.517.0<r<$ 20.5, amounting to 6 978 97769789776\,978\,9776 978 977 stamps. For each source, we also use an RMS map and a PSF image generated from the PSFEx model and oversampled by a factor of 2. This is our ‘parent sample’, which is further characterised in Sect. 3.

### v2

@@
11institutetext: Institute of Physics, Laboratory of Astrophysics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 22institutetext: SCITAS, Ecole Polytechnique Fédérale de Lausanne (EPFL), 1015 Lausanne, Switzerland 33institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí i Franquès 1, 08028 Barcelona, Spain 44institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig de Lluís Companys 23, 08010 Barcelona, Spain 55institutetext: Laboratoire d’Astrophysique de Marseille, UMR7326, Aix-Marseille Université, CNRS, CNES, 13013 Marseille, France 66institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS, and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 77institutetext: Oskar Klein Centre, Department of Physics, Stockholm University, SE-106 91 Stockholm, Sweden 88institutetext: University of Applied Sciences and Arts of Northwestern Switzerland, School of Engineering, 5210 Windisch, Switzerland 99institutetext: Institute of Cosmology and Gravitation, University of Portsmouth, Burnaby Rd, Portsmouth PO1 3FX, UK 1010institutetext: Department of Physics and Astronomy, University of British Columbia, Vancouver, BC V6T 1Z1, Canada 1111institutetext: Canadian Astronomy Data Centre, Herzberg Astronomy and Astrophysics, National Research Council, 5071 West Saanich Rd Victoria BC V9E 2E7, Canada 1212institutetext: LIRA, Observatoire de Paris, Université PSL, CNRS, Place Jules Janssen 92195, Meudon, France 1313institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, ON, N2L 3G1, Canada 1414institutetext: Waterloo Centre for Astrophysics, Waterloo, ON, N2L 3G1, Canada 1515institutetext: Perimeter Institute for Theoretical Physics, 31 Caroline St. N., Waterloo, ON, N2L 2Y5, Canada 1616institutetext: Institute for Astronomy, University of Hawaii, 2680 Woodlawn Drive, Honolulu, HI 96822, USA
@@
J. A. Acevedo Barroso    B. Clément    F. Courbin    R. Gavazzi    C. Lemon    K. Rojas    D. Scott    S. Gwyn    F. Hammer    M. J. Hudson    E. A. Magnier
Recent wide-field galaxy surveys have led to an explosion in the number of galaxy-scale strong gravitational lens candidates. However, the vast majority of them feature massive luminous red galaxies as the main deflectors, with late-type galaxies being vastly under-represented. This work presents a dedicated search for lensing by edge-on late-type galaxies in the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). The search covers $3600\,\deg^{2}$ of $r$-band observations taken from the Canada-France-Hawaii Telescope. We considered all sources with magnitudes in the range $17<r<20.5$, without any colour pre-selection, yielding a parent sample of seven million sources. We characterised our parent sample via the visual inspection of 120 000120\,000 sources selected at random. From it, we estimate, with a 68% confidence interval, that 1 in every 30 00030\,000 sources is an edge-on lens candidate, with at least eight high-quality candidates in the parent sample. This corresponds to one candidate per 17 00017\,000 edge-on late-type galaxies. Our search relied on a convolutional neural network (CNN) to select a reduced sample of candidates, which we followed with a visual inspection to curate the final sample. The CNN was trained from scratch using simulated $r$-band observations of edge-on lenses, and real observations of non-lenses. We found 61 good edge-on lens candidates using the CNN. Moreover, combining the CNN candidates with those found serendipitously and those identified while characterising the parent sample, we discovered 4 grade A, 20 grade B, and 58 grade C edge-on lens candidates, effectively doubling the known sample of these systems. We also discovered 16 grade A, 16 grade B, and 18 grade C lens candidates of other types. Finally, based on the characterisation of the parent sample, we estimate that our search found around 60% of the bright grade A and B edge-on lens candidates within the parent sample.
@@ 1 Introduction
Strong gravitational lensing has become an essential tool for probing the extragalactic Universe. Its applications range from the detection and study of high-redshift sources (e.g., Atek et al. 2023; Bradley et al. 2023), including the mapping of the mass distributions responsible for the lensing effect (e.g., Shajib et al. 2019; Etherington et al. 2022; Mahler et al. 2023), to the constraining of cosmological parameters, most notably $H_{0}$ when the lensed source exhibits time variability (e.g., Wong et al. 2020; Birrer et al. 2020; Kelly et al. 2023; Chen et al. 2024). Additionally, the analysis of flux ratios between multiply lensed images, alongside small-scale distortions in the lensed arcs, allows for the detection of mass clumps in the line of sight (e.g., Vegetti et al. 2010; Nightingale et al. 2023; Nierenberg et al. 2024; Gilman et al. 2024; Nightingale et al. 2024). Furthermore, combining lens modelling with a second probe of mass can break degeneracies in mass models that limit single-probe analyses, such as the baryonic and dark matter degeneracy (e.g., Auger et al. 2010; Schuldt et al. 2019; Sheu et al. 2025), the bulge-disc degeneracy for late-type galaxies (e.g., Dutton et al. 2011; Suyu et al. 2012), or the mass-sheet degeneracy, which affects lens models in general (e.g., Shajib et al. 2023; Khadka et al. 2024). In particular, galaxy-scale lensing, when combined with deep spectroscopic observations or multi-band photometry, can probe the stellar initial mass function (IMF) of the lensing galaxies (see, Ferreras et al. 2010; Czoske et al. 2012; Dutton and Treu 2014; Brewer et al. 2014; Sonnenfeld et al. 2019).
@@
However, these studies are limited by their sample size, and their results can change with improved samples and statistics. This is especially the case for late-type galaxies, for which the number of confirmed lenses amounts to a few dozen objects. Whether the IMF is the same for late-type galaxies as it is for early-type galaxies, in what way the IMF might differ in the bulge from the disc for a given galaxy, and the universality of the IMF for late-type galaxies, are all open questions that can be tackled with large samples of late-type lens galaxies. In this context, the Sloan WFC Edge-on Late-type Lens Survey (SWELLS; WFC – wide field camera) launched a campaign in the early 2010s to find and study lensing by edge-on late-type galaxies (edge-on lenses) while trying to answer the aforementioned questions. They produced a series of six articles covering the curation and characterisation of a spectroscopically selected sample of edge-on lenses (Treu et al. 2011); the combination of lens modelling with spatially resolved kinematics to break the disc-halo degeneracy and probing of the IMF for a single galaxy (Dutton et al. 2011; Barnabè et al. 2012); the systematic comparison between the total mass inside the Einstein radius and the stellar mass estimated from stellar population synthesis (SPS) models to study the IMF for medium-sized samples ( $\approx\,20$) of edge-on lenses (Brewer et al. 2012, 2014); and the study of the IMF normalisation of the bulge and the disc independently for a sample of five edge-on lenses using lens modelling, kinematics, and SPS (Dutton et al. 2013). To date, the SWELLS papers represent the largest study of edge-on lenses, but it is still limited to a few dozen systems.
@@
Galaxy-scale strong lensing events are very rare. The likelihood of observing such events depends on the mass distribution of the deflectors and the redshift distributions of foreground lenses and background sources. As a result, only about one source in thousands exhibits lensed images (e.g., Oguri and Marshall 2010; Collett 2015; Ferrami and Wyithe 2024; Acevedo Barroso et al. 2025), and these sources are not easy to see in ground-based imaging, because of the small image separation on the plane of the sky and because of the low luminosity contrast between the lens light and the background source. This is true when the lens galaxy is a massive luminous red galaxy (LRG), but the situation is even worse for the less massive late-type galaxies. In fact, only about a hundred lens candidates have a late-type deflector, out of approximately twenty thousand lens candidates known by the community (see Vernardos et al., in prep.). Moreover, even though the number of lens candidates is growing orders of magnitude faster since the adoption of machine learning in lens finding, the number of late-type candidates has been growing at a much more modest rate, hinting at the necessity of lens searches that target these objects directly.
@@
Recent massive optical lens searches have targeted wide-sky surveys to maximise the volume probed while also using neural networks to select samples of potential lens candidates. These surveys include the Dark Energy Survey (DES; Jacobs et al. 2019a, b; Rojas et al. 2022; O’Donnell et al. 2022), the Kilo-Degree Survey (KiDS; Petrillo et al. 2017, 2019; He et al. 2020; Li et al. 2020, 2021), the Subaru Hyper Suprime-Cam (HSC; Cañameras et al. 2021; Shu et al. 2022; Jaelani et al. 2024; Schuldt et al. 2025), the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Survey (Huang et al. 2020, 2021; Stein et al. 2022; Storfer et al. 2024), the Euclid Early Release Observations (Nagam et al. 2025), and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Savary et al. 2022). Most of the analyses rely on convolutional neural networks as the main classifier, but visual transformers are also being explored as an alternative (Grespan et al. 2024; Gonzalez et al. 2025).
@@
Given the extremely low prevalence of lensing events, it is common to restrict lens searches to the most massive galaxies, often making a colour pre-selection for LRG. These pre-selections remove most of the late-type galaxy lenses, excluding them from the lens search even before the CNN inspect them. However, even if there is no colour pre-selection, there are still not enough confirmed lenses to train neural networks using only real observations. Instead, the CNN are trained on mock observations simulating the lensing effect. These mock lenses are being refined, from the first ones by Petrillo et al. (2017) randomly adding fully simulated lensed sources to LRG to the recent use of real deep HST observations for the potential sources and the dynamics of the potential deflectors to select a reasonable lensing mass (Cañameras et al. 2020; Rojas et al. 2022; Savary et al. 2022). The problem with this methodology is that it limits the search to the lens candidates that look the most similar to the mock lenses used to train the networks. And since these studies mostly use LRG for the deflector galaxies, the late-type lenses are also indirectly excluded, even if they are part of the parent sample inspected by the networks.
@@
In this work, we search for galaxy-scale lensing by late-type edge-on disc galaxies (hereafter ‘edge-on lenses’) in the Canada-France Imaging Survey (CFIS), now a part of the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). We chose this survey because of its excellent image quality, large area coverage, and its complete overlap with the on-going Euclid Wide Survey (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025). This work represents the second lens search specifically targeting edge-on lenses in optical data, after Sygnet et al. (2010), who also used data from the Canada-France-Hawaii Telescope (CFHT). However, this is the first time it is being done using machine learning and the methodology popularised in the past decade. Our main objective is to substantially increase the sample of known edge-on lens candidates, while also exploring why they are not being found by the recent massive lens searches. Additionally, we aim to estimate the prevalence and total number of edge-on lenses in the data and to develop a framework to evaluate the performance of our CNN search without relying on simulated lenses.
This paper is organised as follows. We present the UNIONS-CFIS observations and the selection of the parent sample in Sect. 2. Then, in Sect. 3, we characterise the parent sample by estimating the number of lenses in it. We describe the process to simulate mock observations of edge-on lenses in Sect. 4, and present the training and deployment of the CNN in Sect. 5. We present spectroscopic follow-up of two candidates from the Nordic Optical Telescope (NOT) in Sect. 6. Finally, we discuss our findings and conclude in Tables 2 and 8.
@@ 2 Data
We used r-band data from CFIS data release 3, which was conducted using the 3.6-m CFHT on Maunakea, Hawaii. CFIS provides deep r- and u-band imaging to UNIONS, a collaboration of wide-field imaging surveys of the northern sky involving CFIS, as well as members of the Panoramic Survey Telescope And Rapid Response System (Pan-STARRS) team, the Waterloo Hawaii Institute for Astronomy G-band Survey (WHIGS) team, and the Wide Imaging with Subaru Hyper Suprime-Cam of the Euclid Sky (WISHES) team. Deep i-band and moderately deep z-band imaging data are provided by Pan-STARRS, deep g-band data are provided by WHIGS, and deep z-band data are provided by the WISHES team. UNIONS aims to provide ground-based complementary data for the Euclid mission, as well as maximising the science output of these ground-based large-area deep imaging surveys (Ibata et al. 2017; Guinot et al. 2022).
The CFIS data release 3 covers $7000$\,\rm{deg}^{2}$ down to a median depth of $24.3$ in the u-band, and $3600$\,\rm{deg}^{2}$ down to a median depth of $25.2$ in the r-band. Both depths are for point sources with $5\,\sigma$ detection in a 2 aperture. The median seeing of the data is 0.8 and 0.6 for the u- and r-bands, respectively. The individual observations were reduced, calibrated and co-added at the Canadian Astronomy Data Centre (CADC) using a modified version of the MegaPipe pipeline (Gwyn 2008, 2019). The final images have a pixel size of 0.1857. Additionally, models of the point spread function (PSF) and its spatial variations were produced for every co-added frame using PSFEx (Bertin 2011). In the present work, we focus on CFIS r-band data due to their exquisite seeing quality. This is crucial for our specific search, since late-type edge-on galaxies have a smaller lensing cross-section than the typical LRG used in other searches. As a consequence, the potential lensed images are more often hidden in the glare of the lens light than for systems with early-type lens galaxies.